Operation Summary for October 1, 2010 - October 4, 2010 |
The Integrated Luminosity for the period from 9/27/10 to 10/4/10 was 62.73 pb-1. Friday Oct 1 The day shift began with Operations monitoring Tevatron (TeV) store 8137 (237.06E30), with no Recycler stash or Antiproton Source (Pbar) stack due to the Main Injector (MI) and Booster accesses, with Linac personnel conducting maintenance, with the NuMI experiment off for target work, with MiniBooNE off due to the Booster access, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding store. During the day shift, Operations established beam to the Booster dump at 2 PM. Operations reported that the MI turn-on was delayed due to communications issues with the MI-50 power supply PLC; EE Support personnel investigated. Operations established beam to Pbar at 3 PM. During the evening shift, around 10:39 PM Operations reported that store 8137 had aborted when the power supply in TeV CAMAC crate $0B failed. Operators changed out the power supply. Operations put the TeV into HEP shot setup at 11:45 PM. Saturday Oct 2 The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (236.1E10), with Pbar stacking (22.6E10), and with MiniBooNE taking beam. During the midnight shift, Operations reported that problems with the TeV sector D1 BPM house and a Booster RF station (BRF6) delayed the loading of final protons and shot setup by ~1.5 hours. After operators recovered BRF6, shot setup continued. Operations established store 8138 at 2:45 AM with an average initial luminosity (AIL) of 229E30. During the day shift, Operations reported that the MI power supply interlock loop dropped out 3 times when the MI-10U VCB opened without any obvious trip status. The VCB closed in each time and operators reestablished the MI ramp. LRF3 tripped numerous times during the afternoon. Operations called in Linac personnel to investigate the LRF3 weak power amplifier. During the evening shift, Operations terminated store 8138 at 4:40 PM, put the TeV into shot setup, and established store 8139 at 5:45 PM with an AIL of 305E30. Sunday Oct 3 The midnight shift began with Operations monitoring store 8139 and the stash, with Pbar stacking, and with MiniBooNE taking beam. During the midnight shift, Operations terminated store 8139 at 6:45 AM, put the TeV into shot setup, and established store 8140 at 8 AM with an AIL of 290E30. During the day shift, Operations reported that the MINOS sump pump #2 tripped off. Operations contacted FESS personnel to investigate. During the evening shift, Operations reported that a beam permit card for the Booster failed. Operations called in Interlock personnel and they reported finding and fixing a faulted MUX signal cable. Operations reestablished Booster beam at 9:40 PM. Operations terminated store 8140 around 11:41 PM and put the TeV into shot setup at 11:53 PM. Monday Oct 4 The midnight shift began with the TeV in shot setup, with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE taking beam. During the midnight shift, Operations established store 8141 at 1:05 AM with an AIL of 287E30. The Plans for Monday The plans for today are to stack, stash, and store. The Run Coordinator will allow MI personnel to tune coalescing and for Switchyard personnel to conduct calibration runs. |
|
More Information Comments and Suggestions |
| last modified 10/04/2010 email Fermilab |
| Security, Privacy, Legal | |